# Semi-decentralized Training of Spatio-Temporal Graph Neural Networks for Traffic Prediction

### v1

@@ Abstract
In smart mobility, large networks of geographically distributed sensors produce vast amounts of high-frequency spatio-temporal data that must be processed in real time to avoid major disruptions. Traditional centralized approaches are increasingly unsuitable to this task, as they struggle to scale with expanding sensor networks, and reliability issues in central components can easily affect the whole deployment. To address these challenges, we explore and adapt semi-decentralized training techniques for Spatio-Temporal Graph Neural Networks (ST-GNNs) in smart mobility domain. We implement a simulation framework where sensors are grouped by proximity into multiple cloudlets, each handling a subgraph of the traffic graph, fetching node features from other cloudlets to train its own local ST-GNN model, and exchanging model updates with other cloudlets to ensure consistency, enhancing scalability and removing reliance on a centralized aggregator. We perform extensive comparative evaluation of four different ST-GNN training setups—centralized, traditional FL, server-free FL, and Gossip Learning—on large-scale traffic datasets, the METR-LA and PeMS-BAY datasets, for short-, mid-, and long-term vehicle speed predictions. Experimental results show that semi-decentralized setups are comparable to centralized approaches in performance metrics, while offering advantages in terms of scalability and fault tolerance. In addition, we highlight often overlooked issues in existing literature for distributed ST-GNNs, such as the variation in model performance across different geographical areas due to region-specific traffic patterns, and the significant communication overhead and computational costs that arise from the large receptive field of GNNs, leading to substantial data transfers and increased computation of partial embeddings.
@@ I Introduction
Like other prediction tasks, traffic forecasting methods can generally be divided into classical statistical methods like autoregressive integrated moving average (ARIMA) [7] and Deep Learning (DL)-based methods such as Long Short-Term Memory (LSTM) [8] and Diffusion Convolutional Recurrent Neural Networks (DCRNN) [9]. However, both approaches only use temporal dependencies, overlooking spatial ones.
Graph Neural Networks (GNNs) [10] address these issues by modeling traffic data as a graph, where nodes represents a physical location, and edges spatial connections between these locations. Specifically, Spatio-Temporal Graph Neural Networks (ST-GNNs) [11] extend GNNs to capture both spatial and temporal dependencies, making them more effective than other DL-based approaches for traffic forecasting. Thus, numerous ST-GNN architectures have been developed and tested for this task [11].
However, most previous works have focus on ST-GNN trained in centralized environments, where data from all sensors is continuously collected. But the task of collecting data, performing DL training and inference, and sending back commands is challenging to perform reliably in real-time in a centralized system. As sensors networks expand, scaling a central control system to handle the increased volume of data in real-time with low-latency becomes increasingly difficult. Furthermore, any issues in the central control system can easily result in a complete interruption of the service, requiring significant investments to maximize reliability. Given the key role played by mobility infrastructure in our societies, a centralized system can also become a significant target for hostile actors and a major cyber-weakness in strategic infrastructure management.
@@
  • To our knowledge, we are the first to investigate semi-decentralize ST-GNN training with geographically distributed cloudlets
@@
  • Using our framework, we implement and compare four different training approaches—centralized, traditional FL, server-free FL, and Gossip Learning—providing the first comprehensive analysis of the benefits and drawback of these approaches for our targeted use-case, assessing their suitability for real-world deployment and informing future research endeavours.
  • We evaluate these approaches on two real-world datasets, METR-LA and PeMS-BAY [12], across short-, mid-, and long-term traffic predictions, with a detailed analysis of model performance, cloudlet-specific metrics, communication overhead and computational costs.
Our experiments show that semi-decentralized approaches showed competitive performance against the centralized approach, especially for short-term prediction. Crucially, our analysis highlights the variation in model performance across different geographical areas due to region-specific traffic patterns, and the significant communication overhead and computational costs that arise from the large receptive field of GNNs, leading to substantial data transfers and increased computation of partial embeddings. This insight sheds light on a notable challenges in distributed ST-GNN training, which to our knowledge has not been explored in existing literature.
@@ II-A Traffic prediction
Traffic prediction is one of the most common and critical time-series forecasting tasks, as it estimates traffic metrics, such as speed, volume, and density to monitor current traffic conditions and predict future trends. It is generally classified into three forecasting horizons: short-term (15 min), mid-term (30 min), and long-term (60 min). In the context of ST-GNNs, the traffic system is represented as a spatial graph at the t-th time step, denoted by $\mathcal{G}_{t}$ = ($\mathcal{V}_{t}$, $\mathcal{E}$, W), where $v_{i},v_{j}\in\mathcal{V}_{t}$ is a finite set of nodes corresponding to observations from n IoT-devices in the traffic network, $(v_{i},v_{j})\in\mathcal{E}$ is a set of edges connecting the IoT-devices, and $W\in\mathbb{R}^{n\times n}$ is a weighted adjacency matrix that encodes the spatial relationships between these IoT-devices. The spatial component is defined by the physical locations of IoT devices, while the temporal component arises from traffic features that evolve over time, influencing the graph’s dynamics. While this work focuses on traffic prediction, methods for other smart mobility tasks can be generalized to this domain due to their shared reliance on spatio-temporal data.
@@ III-A Scenario
The traffic prediction scenario involves IoT-devices (i.e., sensors) deployed along highways to monitor real-time traffic conditions such as vehicle speed, volume, and density at fixed time intervals. To streamline data processing, IoT-devices are assigned to local base stations (BS), referred to as cloudlets, based on their physical proximity and communication range. The communication between IoT-devices and its BS is facilitated using Low-Power Wide-Area Network (LPWAN) protocols, such as LoRa or NB-IoT, which enable efficient, long-range, and low-power data transmission. Base stations, in addition to serving as communication hubs, are equipped with moderate computational resources. This existing computational power can be leveraged for training ML models, such as ST-GNNs.
@@ IV-A Datasets
The spatial adjacency matrix is constructed from the actual road network based on distance using the formula from ChebNet [20]. Specifically, it encodes the spatial relationships between nodes by assigning weights to edges based on the inverse of the pairwise distances between connected nodes, effectively capturing the connectivity and proximity of locations in the traffic network. Additionally, the data were preprocessed for efficient model training and evaluation. For example, if we want to utilize the historical data spanning one hour to predict the data the next hour, we pack the sequence data in group of 12 and convert it into an instance. The dataset was then split into training, validation, and test sets in a 70:15:15 ratio. Additionally, to mitigate the impact of varying data magnitudes, we applied standardization to normalize the original data values, reducing the impact of the large difference in value.
@@ IV-B Evaluation Metrics
In our experiments, we evaluate the performance of different setups using Weighted Mean Absolute Percentage Error (WMAPE), alongside the widely adopted Mean Absolute Error (MAE) and Root Mean Square Error (RMSE). We omit the formulas for MAE and RMSE due to their wide adoption. The WMAPE formula is:
$\text{WMAPE}(\bm{x},\bm{\hat{x}})=\frac{\sum_{i}{\mid x_{i}-\hat{x}_{i}\mid}}{\sum_{i}{\hat{x}_{i}}}*100\%$ (1)
@@
We compute these metrics across three forecasting horizons—–short-, mid-, and long-term—to capture performance variations over time. For centralized and FL setups, metrics are derived from the main server’s aggregated model, while for server-free FL and Gossip Learning, they are computed as a weighted average of individual cloudlet predictions. This ensures a consistent and fair comparison across all approaches.
Figure 2: Sensor partitioning to cloudlets based on communication range
@@ IV-C Experimental Setup
TABLE II: Performance Comparison (MAE [$mile/h$]/RMSE [$mile^{2}/h^{2}$]/WMAPE [%]) of four different setups
@@
For model architecture, we used 2 spatio-temporal blocks (ST-blocks), with GLU activation function, a learning rate of 0.0001, dropout of 0.5, weight decay of 0.00001, and a batch size of 32. MAE is used as the training loss, Adam was chosen as the optimization algorithm, and learning rate scheduler was applied with the StepLR strategy, where the step size was set to 5, and the decay factor was set to 0.7. The temporal and spatial kernel size was set to 3 for all experiments, with Chebyshev convolution as convolutional layer. All experiments use 60 minutes as the historical time window, i.e., 12 observed data points.
Additionally, sensors were distributed across 7 cloudlets based on proximity and communication range. Each cloudlet can communicate with other cloudlets and IoT-devices only if they are within an 8 km range. This limitation directly impacts the server-free FL approach by restricting the number of cloudlets that can exchange model updates. In contrast, this constraint does not affect Gossip Learning, as model updates are sent to a randomly selected cloudlet across the entire network, regardless of proximity. Fig. 2 shows how sensors are assigned to cloudlets and their respective communication ranges for both the METR-LA and PeMS-BAY datasets. This partitioning reflects the semi-decentralized structure used for our distributed training approach.
@@ V-B Cloudlet Metric Analysis
Additionally, while global weighted averages reported in Table II provide a summary of performance, they fail to capture the significant differences in performance across individual cloudlets. Previous works have primarily focused on global averages, overlooking notable challenge of heterogeneous performance among cloudlets. This oversight can lead to an incomplete understanding of model performance in distributed traffic forecasting systems, resulting in unexpected discrepancies, suboptimal decision-making and potential disruptions in real-world usage compared to testing.
@@
TABLE III: FLOPs and average model and node feature transer size per cloudlet
@@
METR-LA Centralized - 0.24T - 0.68
Traditional FL 0.13 1.56T 0.16M 3.69
Server-free FL 0.23 1.56T 0.52M 3.69
Gossip Learning 0.13 1.56T 0.42M 3.69
PeMS-BAY Centralized - 0.58T - 1.61
Traditional FL 0.13 3.89T 0.16M 9.44
Server-free FL 0.26 3.89T 0.56M 9.44
Gossip Learning 0.13 3.89T 0.42M 9.44
@@ V-C Analysis of semi-decentralized overheads
In contrast, model transfer per epoch is relatively small due to the compact size of ST-GCN models. While model transfers accumulate over time, contributing to total communication overhead, the per-epoch cost remains in the order of megabytes. As shown in Figure 4, distributed setups, particularly server-free FL, require more epochs to converge than centralized training, increasing total model transfer costs. Despite this, even in server-free FL’s worst-case scenario, where all cloudlets communicate with each other, model transfer remains a minor overhead.
@@ VI Related Work
Training and testing GNN models on large graphs demand substantial memory and processing resources due to the interdependence of graph nodes [21]. This creates a scalability and reliability issues for centralized GNN operations [22], especially when dealing with large-scale graphs or sensitive data collected from distributed sensors in real-time traffic prediction.
@@ VI-B Traditional Federated Learning ST-GNN training
Several works have applied traditional FL to train ST-GNN models for traffic prediction [24, 25, 26, 27]. One of the primary advantages of traditional FL is enhanced scalability. By distributing the training process, it leverages the computational resources of multiple devices, which allows larger datasets to be processed without overloading any single server. While this approach improves on centralized training, it still relies on a central aggregator, meaning that it shares the same issues as centralized approach, as was already described above.
@@ VI-C Decentralized and semi-decentralized ST-GNN training
One relevant work by L. Giaretta et. al. [28] explores the challenges and solutions associated with fully decentralized training. However, their solution is not tailored to ST-GNNs, but rather to Graph Convolutional Network (GCN), limiting its applicability to traffic prediction tasks that require both spatial and temporal correlations.

### v2

@@ Abstract
In smart mobility, large networks of geographically distributed sensors produce vast amounts of high-frequency spatio-temporal data that must be processed in real time to avoid major disruptions. Traditional centralized approaches are increasingly unsuitable to this task, as they struggle to scale with expanding sensor networks, and reliability issues in central components can easily affect the whole deployment. To address these challenges, we explore and adapt semi-decentralized training techniques for Spatio-Temporal Graph Neural Networks (ST-GNNs) in the smart mobility domain. We implement a simulation framework where sensors are grouped by proximity into multiple cloudlets, each handling a subgraph of the traffic graph, fetching node features from other cloudlets to train its own local ST-GNN model, and exchanging model updates with other cloudlets to ensure consistency, enhancing scalability and removing reliance on a centralized aggregator. We perform extensive comparative evaluation of four different ST-GNN training setups—centralized, traditional FL, server-free FL, and Gossip Learning—on large-scale traffic datasets, the METR-LA and PeMS-BAY datasets, for short-, mid-, and long-term vehicle speed predictions. Experimental results show that semi-decentralized setups are comparable to centralized approaches in performance metrics, while offering advantages in terms of scalability and fault tolerance. In addition, we highlight often overlooked issues in existing literature for distributed ST-GNNs, such as the variation in model performance across different geographical areas due to region-specific traffic patterns, and the significant communication overhead and computational costs. However, due to the planar nature of graphs, per-cloudlet costs remain consistent as the network grows, unlike the growing costs in a centralized approach.
@@ I Introduction
Like other prediction tasks, traffic forecasting methods can generally be divided into classical statistical methods like autoregressive integrated moving average (ARIMA) [7] and Machine Learning (ML) methods. The latter can be further split into “traditional” ML methods, such as gradient boosting [8], and Deep Learning (DL)-based methods including Long Short-Term Memory (LSTM) [9] and Diffusion Convolutional Recurrent Neural Networks (DCRNN) [10]. However, all of these methods only exploit temporal dependencies, overlooking spatial ones.
Graph Neural Networks (GNNs) [11] address these issues by modeling traffic data as a graph, where nodes represent physical locations, and edges spatial connections between these locations. Specifically, Spatio-Temporal Graph Neural Networks (ST-GNNs) [12] extend GNNs to capture both spatial and temporal dependencies, making them more effective than other DL-based approaches for traffic forecasting. Thus, numerous ST-GNN architectures have been developed and tested for this task [12].
However, most previous works have focused on ST-GNN trained in centralized environments, where data from all sensors is continuously collected. But the task of collecting data, performing DL training and inference, and sending back commands is challenging to perform reliably in real-time in a centralized system. As sensor networks expand, scaling a central control system to handle the increased volume of data in real-time with low latency becomes increasingly difficult. Furthermore, any issues in the central control system can easily result in a complete interruption of the service, requiring significant investments to maximize reliability. Given the key role played by mobility infrastructure in our societies, a centralized system can also become a significant target for hostile actors and a major cyber-weakness in strategic infrastructure management.
@@
  • To our knowledge, we are the first to investigate semi-decentralized ST-GNN training with geographically distributed cloudlets
@@
  • Using our framework, we implement and compare four different training approaches—centralized, traditional FL, server-free FL, and Gossip Learning—providing the first comprehensive analysis of the benefits and drawbacks of these approaches for our targeted use-case, assessing their suitability for real-world deployment and informing future research endeavours.
  • We evaluate these approaches on two real-world datasets, METR-LA and PeMS-BAY [13], across short-, mid-, and long-term traffic predictions, with a detailed analysis of model performance, cloudlet-specific metrics, communication overhead, and computational costs.
Our experiments show that semi-decentralized approaches perform competitively against the centralized approach, especially for short-term prediction. Crucially, we provide insights on notable challenges in distributed ST-GNN training, which to our knowledge has not been explored in existing literature, such as highlighting the variation in model performance across different geographical areas due to region-specific traffic patterns, and the significant communication overhead and computational costs that arise from the large receptive field of GNNs, leading to substantial data transfers and increased computation of partial embeddings. However, communication overhead is mitigated by the planar structure of graphs, keeping per-cloudlet costs consistent as the network grows.
@@ II-A Traffic prediction
Traffic prediction is one of the most common and critical time-series forecasting tasks, as it estimates traffic metrics, such as speed, volume, and density to monitor current traffic conditions and predict future trends. It is generally classified into three forecasting horizons: short-term (15 min), mid-term (30 min), and long-term (60 min) [14]. In the context of ST-GNNs, the traffic system is represented as a spatial graph at the t-th time step, denoted by $\mathcal{G}_{t}$ = ($\mathcal{V}_{t}$, $\mathcal{E}$, W), where $v_{i},v_{j}\in\mathcal{V}_{t}$ is a finite set of nodes corresponding to observations from n IoT-devices in the traffic network, $(v_{i},v_{j})\in\mathcal{E}$ is a set of edges connecting the IoT-devices, and $W\in\mathbb{R}^{n\times n}$ is a weighted adjacency matrix that encodes the spatial relationships between these IoT-devices. The spatial component is defined by the physical locations of IoT devices, while the temporal component arises from traffic features that evolve over time, influencing the graph’s dynamics. While this work focuses on traffic prediction, methods for other smart mobility tasks can be generalized to this domain due to their shared reliance on spatio-temporal data.
@@ II-E Gossip Learning
II-F Distributed training
In distributed machine learning, training paradigms can be categorized based on their system architecture and coordination mechanisms:
  • Centralized Training: Data is collected and processed on a single server or coordinator, which performs model updates.
  • Semi-Decentralized Training: A hybrid approach where devices are organized into clusters or hierarchies. Local coordination occurs within clusters, while limited global coordination ensures model consistency.
  • Fully-Decentralized Training: Devices collaborate in a peer-to-peer manner, exchanging model updates directly without relying on a central coordinator.
@@ III-A Scenario
The traffic prediction scenario involves IoT-devices (i.e., sensors) deployed along highways at fixed locations to capture traffic metrics in real-time at fixed time intervals. To streamline data processing, IoT-devices are assigned to local base stations (BS) at fixed locations, referred to as cloudlets, based on their physical proximity and communication range. The communication between IoT-devices and its BS is facilitated using Low-Power Wide-Area Network (LPWAN) protocols, such as LoRa or NB-IoT, which enable efficient, long-range, and low-power data transmission. Base stations, in addition to serving as communication hubs, are equipped with moderate computational resources. This existing computational power can be leveraged for training ML models, such as ST-GNNs.
@@ IV-A Datasets
The spatial adjacency matrix is constructed from the actual road network based on distance using the formula from ChebNet [22]. Specifically, it encodes the spatial relationships between nodes by assigning weights to edges based on the inverse of the pairwise distances between connected nodes, effectively capturing the connectivity and proximity of locations in the traffic network. Additionally, the data were preprocessed for efficient model training and evaluation. For example, if we want to utilize the historical data spanning one hour to predict the data the next hour, we pack the sequence data in group of 12 and convert it into an instance. The dataset was then split into training, validation, and test sets in a 70:15:15 ratio. The validation set is used for early stopping and to select the best model during training. The final results are then reported on the test dataset using the best-performing model identified through validation. Additionally, to mitigate the impact of varying data magnitudes, we applied standardization to normalize the original data values, reducing the impact of the large difference in value.
@@ IV-B Evaluation Metrics
In our experiments, we evaluate the performance of different setups using Weighted Mean Absolute Percentage Error (WMAPE), in addition to the widely adopted Mean Absolute Error (MAE) and Root Mean Square Error (RMSE). All metrics are computed after rescaling the predictions back to the original data range, before standardization. We use WMAPE because it accounts for the varying magnitudes of traffic speeds across regions, providing a more balanced evaluation, compared to Mean Absolute Percentage Error (MAPE). We omit the formulas for MAE and RMSE due to their wide adoption. The WMAPE formula is as follows:
$\text{WMAPE}(\bm{x},\bm{\hat{x}})=\frac{\sum_{i}{\mid x_{i}-\hat{x}_{i}\mid}}{\sum_{i}{\hat{x}_{i}}}*100\%,$ (1)
@@
We compute these metrics across three forecasting horizons—–short-, mid-, and long-term—to capture fluctuations in performance over time. For centralized and FL setups, the metrics are derived from the main server’s aggregated model, while for server-free FL and Gossip Learning, they are computed as a weighted average of the individual cloudlet predictions. This ensures a consistent and fair comparison across all approaches.
Figure 2: Sensor assignment to cloudlets based on communication range
@@ IV-C Experimental Setup
TABLE II: Performance Comparison (MAE [$mile/h$]/RMSE [$mile/h$]/WMAPE [%]) of four different setups
@@
For model architecture, we used 2 spatio-temporal blocks (ST-blocks), with GLU activation function, a learning rate of 0.0001, dropout of 0.5, weight decay of 0.00001, and a batch size of 32. MAE is used as the training loss, Adam was chosen as the optimization algorithm, and learning rate scheduler was applied with the StepLR strategy, where the step size was set to 5, and the decay factor was set to 0.7. The temporal and spatial kernel size was set to 3 for all experiments, with Chebyshev convolution as convolutional layer. All experiments use 60 minutes as the historical time window, i.e., 12 observed data points. The hyperparameters used in our experiments are based on previously established configurations that have been optimized for the given model architecture and datasets [16]. We adopt these settings to ensure comparability with prior work and maintain consistency in evaluating different training approaches.
Additionally, sensors were distributed across 7 cloudlets based on proximity and communication range. While algorithms such as k-means clustering could have been used to distribute sensors to cloudlets, we manually defined cloudlet placement to ensure full coverage of the geographical area in each dataset. Each cloudlet can communicate with other cloudlets and with IoT devices only if they are within an 8 km range. This limitation directly impacts the server-free FL approach by restricting the number of cloudlets that can exchange model updates. In contrast, this constraint does not affect Gossip Learning, as model updates are sent to a randomly selected cloudlet across the entire network, regardless of proximity. Fig. 2 shows how sensors are assigned to cloudlets and their respective communication ranges for both the METR-LA and PeMS-BAY datasets. This partitioning reflects the semi-decentralized structure used for our distributed training approach.
Our experimental setup is designed to analyze global average prediction performance of each training approach using evaluation metrics, while also providing a comprehensive view of how errors distribute geographically across cloudlets for different prediction horizons. Additionally, we analyze the communication and computational overhead associated with each setup, with particular attention to the high cost of node feature transfer in high-density graphs.
@@ V-B Cloudlet Metric Analysis
Additionally, while global weighted averages reported in Table II provide a summary of performance, they fail to capture the significant differences in performance across individual cloudlets. Previous works have primarily focused on global averages, overlooking the notable challenge of heterogeneous performance among cloudlets. This oversight can lead to an incomplete understanding of model performance in distributed traffic forecasting systems, resulting in unexpected discrepancies, suboptimal decision-making, and potential disruptions in real-world usage compared to testing.
@@
TABLE III: FLOPs and average model and node feature transfer size
@@
METR-LA Centralized - 1.68T - 4.76
Traditional FL 0.91 / 7 10.92T / 7 1.12M / 7 25.83 / 7
Server-free FL 1.61 / 7 10.92T / 7 3.64M / 7 25.83 / 7
Gossip Learning 0.91 / 7 10.92T / 7 2.94M / 7 25.83 / 7
PeMS-BAY Centralized - 4.06T - 11.27
Traditional FL 0.91 / 7 27.23T / 7 1.12M / 7 66.08 / 7
Server-free FL 1.82 / 7 27.23T / 7 3.92M / 7 66.08 / 7
Gossip Learning 0.91 / 7 27.23T / 7 2.94M / 7 66.08 / 7
@@ V-C Analysis of semi-decentralized overheads
In contrast, model transfer per epoch is relatively small due to the compact size of ST-GCN models. While model transfers accumulate over time, contributing to total communication overhead, the per-epoch cost remains in the order of megabytes. As shown in Figure 4, distributed setups, particularly server-free FL, require more epochs to converge than centralized training, increasing total model transfer costs. This slower convergence is further influenced by the high amount of duplicated node features across cloudlets in semi-decentralized approaches. Despite this, even in server-free FL’s worst-case scenario, where all cloudlets communicate with each other, model transfer remains a minor overhead.
@@ VI Related Work
Training and testing GNN models on large graphs demand substantial memory and processing resources due to the interdependence of graph nodes [23]. This creates scalability and reliability issues for centralized GNN operations [24], especially when dealing with large-scale graphs or sensitive data collected from distributed sensors in real-time traffic prediction.
@@ VI-B Traditional Federated Learning ST-GNN training
Several works have applied traditional FL to train ST-GNN models for traffic prediction [26, 27, 28, 29]. One of the primary advantages of traditional FL is enhanced scalability. By distributing the training process, it leverages the computational resources of multiple devices, which allows larger datasets to be processed without overloading any single server. While this approach improves on centralized training, it still relies on a central aggregator, meaning that it shares the same issues as the centralized approach, as was already described above.
@@ VI-C Decentralized and semi-decentralized ST-GNN training
One relevant work by L. Giaretta et. al. [30] explores the challenges and solutions associated with fully decentralized training. However, their solution is not tailored to ST-GNNs, but rather to Graph Convolutional Networks (GCNs), limiting its applicability to traffic prediction tasks that require both spatial and temporal capabilities.
@@ VII-B Cloudlet personalization
VII-C Cloudlet optimization
An important limitation of our current approach is the predefined placement of cloudlets, and communication range, which has not been optimized for coverage or efficiency. The current cloudlet locations were chosen based on a simplified partitioning strategy, ensuring communication between cloudlets, and ensuring that each node is assigned to a nearby cloudlet. Future work should explore strategies for optimizing both the number and placement of cloudlets to enhance coverage and communication efficiency, while maintaining a low performance gap compared to the centralized setup, and while also leveraging the pre-existing deployment of mobile network base-stations to determine practical cloudlet placement, ensuring realistic deployment scenarios.